OpenStudy (vortish):

8 inches of heavy wet snow is the equivalent to 1 inch of rain estimate the water content of 15 inches of heavy wet snow

OpenStudy (mathstudent55):

Use a proportion. heavy wet snow/rain = 8/1 = 15/x Solve 8/1 = 15/x by cross multiplying.

OpenStudy (vortish):